### v1

# DDAD: A Two-pronged Adversarial Defense Based on Distributional Discrepancy

Abstract
Statistical adversarial data detection (SADD) detects whether an upcoming batch contains adversarial examples (AEs) by measuring the distributional discrepancies between clean examples (CEs) and AEs. In this paper, we reveal the potential strength of SADD-based methods by theoretically showing that minimizing distributional discrepancy can help reduce the expected loss on AEs. Nevertheless, despite these advantages, SADD-based methods have a potential limitation: they discard inputs that are detected as AEs, leading to the loss of clean information within those inputs. To address this limitation, we propose a two-pronged adversarial defense method, named Distributional-Discrepancy-based Adversarial Defense (DDAD). In the training phase, DDAD first optimizes the test power of the maximum mean discrepancy (MMD) to derive MMD-OPT, and then trains a denoiser by minimizing the MMD-OPT between CEs and AEs. In the inference phase, DDAD first leverages MMD-OPT to differentiate CEs and AEs, and then applies a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the distributional-discrepancy-based denoiser. Extensive experiments show that DDAD outperforms current state-of-the-art (SOTA) defense methods by notably improving clean and robust accuracy on CIFAR-10 and ImageNet-1K against adaptive white-box attacks.
Machine Learning, ICML
[…]
1 Introduction
The discovery of adversarial examples (AEs) has raised a security concern for artificial intelligence techniques in recent decades (Szegedy et al., 2014; Goodfellow et al., 2015). AEs are often crafted by adding imperceptible noise to clean examples (CEs), which can easily mislead a well-trained deep learning model to make wrong predictions. Considering the extensive use of deep learning systems, AEs pose a significant security threat for real-world applications (Sharif et al., 2016; Dong et al., 2019; Finlayson et al., 2019; Cao et al., 2021; Jing et al., 2021). Therefore, it is imperative to develop advanced defense methods to defend against AEs (Goodfellow et al., 2015; Madry et al., 2018; Zhang et al., 2019; Wang et al., 2020; Yoon et al., 2021; Nie et al., 2022).
Recently, statistical adversarial data detection (SADD) has gained increasing attention due to its effectiveness in detecting AEs (Gao et al., 2021; Zhang et al., 2023). Unlike other detection-based methods that train a detector for specific classifiers (Stutz et al., 2020; Deng et al., 2021; Pang et al., 2022b), SADD leverages statistical methods, for example, maximum mean discrepancy (MMD) (Gretton et al., 2012), to measure the discrepancies between the clean and adversarial distributions. Given the fact that clean and adversarial data are from different distributions, SADD-based methods have been shown empirically effective against adversarial attacks (Gao et al., 2021; Zhang et al., 2023).
To understand the intrinsic strength of SADD-based methods from a theoretical standpoint, we establish a relationship between distributional discrepancy and the expected loss on adversarial data. Our theoretical analysis demonstrates that minimizing distributional discrepancy can help reduce the expected loss on adversarial data, revealing the potential value of leveraging distributional discrepancy to design more effective defense methods (see Section 2 and 3).
However, despite their effectiveness from both empirical and theoretical perspectives, detection-based methods (e.g., SADD-based methods) have a potential limitation: they discard inputs if they are detected as AEs, leading to the loss of clean information (e.g., semantic information) within those inputs. This issue is more prominent in SADD-based methods, where inputs are often processed in batches, potentially resulting in the unintended loss of some CEs along with AEs if a batch contains a mixture of CEs and AEs (Gao et al., 2021; Zhang et al., 2023). Furthermore, in many domains, obtaining large quantities of high-quality data is challenging due to factors such as cost, privacy concerns, or the rarity of specific data, for example, obtaining medical images for rare diseases is challenging (Litjens et al., 2017). As a result, all possible samples with clean information are critical in these data-scarce domains (Gandhar et al., 2024). Therefore, given the effectiveness of SADD-based methods, the above-mentioned challenges naturally lead us to pose the following question: Can we design an adversarial defense method that leverages the effectiveness of SADD-based methods, while at the same time, preserves all the data before feeding them into a classifier?
The answer to this question is affirmative. Motivated by our theoretical analysis, we propose a two-pronged adversarial defense called Distributional-Discrepancy-based Adversarial Defense (DDAD). Specifically, we leverage an advanced MMD statistic (named MMD-OPT) in our pipeline, which is obtained by maximizing the testing power of MMD (see Algorithm 1). MMD-OPT serves two roles: in the training phase of the denoiser (see Algorithm 2), MMD-OPT serves as a ‘guider’ that can help minimize the distributional discrepancies between AEs and CEs. Then, by simultaneously minimizing the cross-entropy loss, we aim to train a denoiser that can minimize the distributional discrepancy towards the direction of making the classification correct; in the inference phase (see Section 4.3), MMD-OPT serves as a ‘detector’ that can help differentiate CEs and AEs. Then, our method applies a two-pronged process: (1) directly feeding the detected CEs into the classifier, and (2) removing noise from the detected AEs by the denoiser through distributional discrepancy minimization. We provide a visual illustration in Figure 1.
Through extensive evaluations on benchmark image datasets such as CIFAR-10 (Krizhevsky et al., 2009) and Imagenet-1K (Deng et al., 2009), we demonstrate the effectiveness of DDAD in Section 5. Compared to current state-of-the-art (SOTA) adversarial defense methods, DDAD can improve clean and robust accuracy by a notable margin against well-designed adaptive white-box attacks (see Section 5.2). Furthermore, experiments show that DDAD can generalize well against unseen transfer attacks (see Section 5.3).
The success of DDAD in adversarial classification takes root in the following aspects: (1) minimizing distributional discrepancies has the potential to reduce the expected loss on AEs; (2) the two-pronged process combines the strengths of SADD-based and denoiser-based methods while also addressing their potential limitations: SADD-based methods can effectively distinguish AEs from CEs but discard the clean information within AEs. In contrast, denoiser-based methods can handle both data without re-training the downstream task model. However, they cannot distinguish AEs and CEs beforehand, which often results in a drop in clean accuracy. Our method, on the other hand, separates CEs and AEs in the inference phase, thereby keeping the accuracy for CEs nearly unaffected. At the same time, AEs can be properly handled by the denoiser; (3) compared to most denoiser-based methods that rely on density estimation (e.g., Nie et al. (2022) and Lee & Kim (2023)), learning distributional discrepancies is a simpler and more feasible task, especially on large-scale datasets such as ImageNet-1K.
2 Problem Setting
In this section, we discuss the problem setting for the adversarial classification in detail.
We formalize our problem for $K$-class classification as follows. We define a domain as a pair consisting of a distribution $\mathcal{D}$ on inputs $\mathcal{X}$ and a labelling function $f:\mathcal{X}\to\{0,1,...,K\}$. Specifically, we consider a clean domain and an adversarial domain. The clean domain is denoted by $\langle\mathcal{D_{C}},f_{\mathcal{C}}\rangle$, and the adversarial domain is denoted by $\langle\mathcal{D_{A}},f_{\mathcal{A}}\rangle$. We define a hypothesis as a function $h:\mathcal{X}\rightarrow\{0,1,...,K\}$ from the hypothesis space $\mathcal{H}$. The probability according to the distribution $\mathcal{D}$ that a hypothesis $h$ disagrees with a labelling function $f$ (which can also be a hypothesis) is the risk:
$R(h,f,\mathcal{D})=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\left[\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))\right],$
where $\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))$ is a loss function that measures the disagreement between $h(\mathbf{x})$ and $f(\mathbf{x})$.
We consider the clean risk of a hypothesis as $R(h,f_{\mathcal{C}},\mathcal{D_{C}})$, and the adversarial risk as $R(h,f_{\mathcal{A}},\mathcal{D_{A}})$. In our problem, adversarial data are generated based on the given clean data. Therefore, $\mathcal{D_{C}}$ is fixed and we use $\mathbb{D}$ to represent a set of valid adversarial distributions such that all possible $\mathcal{D_{A}}\in\mathbb{D}$.
Assumption 1.
For any valid adversarial attack, adversarial data are generated by adding an $\epsilon$-norm-bounded imperceptible perturbation $\epsilon^{\prime}$ to the given clean data without changing its semantic meaning. Assume a valid ground-truth labelling function $f_{\mathcal{A}}$ exists, $f_{\mathcal{A}}$ satisfies the following property:
$\forall\epsilon^{\prime}\ \mbox{s.t.}\ \|\epsilon^{\prime}\|_{p}\leq\epsilon,\quad f_{\mathcal{A}}(\mathbf{x}+\epsilon^{\prime})=f_{\mathcal{A}}(\mathbf{x}),$
where $\epsilon$ is the maximum allowed perturbation budget, and $\|\cdot\|_{p}$ is the threat model’s $\ell_{p}$ norm.
Assumption 2.
Attacks in the adversarial domain will not change the labelling from the clean ground truth, i.e., mathematically:
$\forall\epsilon^{\prime}\ \mbox{s.t.}\ \|\epsilon^{\prime}\|_{p}\leq\epsilon,\quad f_{\mathcal{A}}(\mathbf{x}+\epsilon^{\prime})=f_{\mathcal{C}}(\mathbf{x}),$
where $\epsilon$ is the maximum allowed perturbation budget.
Corollary 1.
If Assumptions 1 and 2 both hold, then we have:
$\forall\mathbf{x}\in\mathcal{X},\quad f_{\mathcal{C}}(\mathbf{x})=f_{\mathcal{A}}(\mathbf{x}).$
Remark 1.
Assumptions 1 and 2 are more like inherent truths here, as attacks should only generate valid examples that abide by the original label (Bartoldson et al., 2024). Therefore, compared to the setting of common domain adaptation problems (Ben-David et al., 2006, 2010), the ground-truth labelling functions for the clean and adversarial domains are equal in our problem setting.
3 Motivation from Theoretical Justification
In this section, we study a toy setting on the relationship between adversarial risk and distributional discrepancy, aiming to shed some light on designing effective adversarial defense methods.
Simplified problem setting. For simplicity, we analyze our problem for binary classification, i.e., a labelling function $f$ is simplified to $f:\mathcal{X}\to\{0,1\}$ and a hypothesis $h\in\mathcal{H}$ is simplified to $h:\mathcal{X}\rightarrow\{0,1\}$. The loss function is simplified to 0-1 loss (i.e., $\mathcal{L}(h(\mathbf{x}),f(\mathbf{x}))=|h(\mathbf{x})-f(\mathbf{x})|$). Otherwise, other settings (e.g., the definition of risks) are the same as described in Section 2.
Definition 1.
For simplicity, we use $L^{1}$ divergence, one of the most distinguishable metrics, as a natural measure of distributional discrepancies between two distributions:
$d_{1}(\mathcal{D},\mathcal{D^{\prime}})=2\sup_{B\in\mathcal{B}}|\Pr_{\mathcal{D}}[B]-\Pr_{\mathcal{D}^{\prime}}[B]|,$
where $\mathcal{B}$ is the set of measurable subsets under $\mathcal{D}$ and $\mathcal{D^{\prime}}$.
Theorem 1.
For a hypothesis $h\in\mathcal{H}$ and a distribution $\mathcal{D_{A}}\in\mathbb{D}$:
$R(h,f_{\mathcal{A}},\mathcal{D_{A}})\leq R(h,f_{\mathcal{C}},\mathcal{D_{C}})+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}}).$
The proof of Theorem 1 can be found in Appendix A.
Definition 2.
The optimal hypothesis that minimizes the clean risk is defined as:
$h_{\mathcal{C}}^{*}=\operatorname*{arg\,min}_{h\in\mathcal{H}}R(h,f_{\mathcal{C}},\mathcal{D_{C}}).$
Significance of distributional discrepancy to adversarial defense. In our problem, we use a practical setting that an attacker aims to attack a well-trained classifier on clean data (i.e., ideally the clean risk is minimized). According to Theorem 1, we have:
$R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})\leq R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}}).$ (1)
Since $h_{\mathcal{C}}^{*}$, $f_{\mathcal{C}}$ and $\mathcal{D_{C}}$ are fixed, $R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})$ is possibly a small constant (according to Definition 2). In our problem, the objective of an attacker can be considered as finding an optimal $\mathcal{D_{A}}\in\mathbb{D}$ that maximizes $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})$. Now, assume we have a detector that leverages the distributional discrepancies to identify AEs. Then, to break the defense, the attacker must generate AEs that could minimize the distributional discrepancies between CEs and AEs (i.e., to mislead the detector to identify AEs as CEs). However, according to Eq. 1, reducing the distributional discrepancy $d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})$ can help reduce adversarial risk $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})$, which violates the objective of adversarial attacks. This intriguing phenomenon helps explain why SADD-based methods are effective against adaptive attacks in practice and inspires the design of our proposed method in this paper (see Section 4).
Comparison with previous studies. Previous studies have attempted to use distributional discrepancy in adversarial defense. For example, at the early stage of AT, Song et al. (2019) propose to treat adversarial attacks as a domain adaptation problem. However, to the best of our knowledge, the relationship between adversarial risk and distributional discrepancy has not been well investigated yet from a theoretical perspective. In previous domain adaptation literature, the upper bound of the risk on the target domain is always bounded by one extra constant (Mansour et al., 2009; Ben-David et al., 2010), e.g., $R(h_{\mathcal{C}}^{*},f_{\mathcal{A}},\mathcal{D_{A}})\leq R(h_{\mathcal{C}}^{*},f_{\mathcal{C}},\mathcal{D_{C}})+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})+C$. This constant $C$ may prevent decreasing the risk on the target domain from minimizing the distributional discrepancy between the source domain and the target domain. By contrast, we treat adversarial classification as a special domain adaptation problem where the ground truth labelling functions are equivalent for both source and target domain. Based on this, we derive an upper bound without any extra constant, i.e., distributional discrepancy minimization can help reduce the expected loss on adversarial domain.
4 Methodology
Figure 1: The illustration of Distributional-Discrepancy-based Adversarial Defense (DDAD). In the training phase, DDAD first optimizes the test power of the maximum mean discrepancy (MMD) to derive MMD-OPT and then trains a denoiser by minimizing the MMD-OPT between CEs and AEs. Then, by simultaneously minimizing the cross-entropy loss, we aim to obtain a denoiser that can minimize the distributional discrepancy towards the direction of making the classification correct. In the inference phase, DDAD uses MMD-OPT to detect AEs and then denoises them instead of discarding them. Conversely, our method will directly feed detected CEs into the classifier.
Motivated by our theoretical analysis in Section 3, we propose a two-pronged adversarial defense method called Distributional-Discrepancy-based Adversarial Defense (DDAD). In this section, we will first introduce the concepts of maximum mean discrepancy (MMD). This will be followed by a detailed discussion of the training and inference process of DDAD. We provide a visual illustration for DDAD in Figure 1 and a detailed description of mathematical notations in Appendix B.
4.1 Preliminary
Maximum mean discrepancy. The problem of using $L^{1}$ divergence in practice is that it does not any unbiased estimators. This is because the supremum can hardly be approximated by finite samples. Hence, in practice, it is challenging to estimate $L^{1}$ divergence accurately, especially in high-dimensional settings, where the bias and variance of the estimation can become significant. Therefore, in this paper, we use MMD to measure the distributional discrepancies between AEs and CEs. MMD has unbiased estimators and can effectively distinguish the difference between two distributions using small batches of data (Liu et al., 2020; Gao et al., 2021; Zhang et al., 2023). Following Gretton et al. (2012), let $\mathcal{X}\subset\mathbb{R}^{d}$ denote a separable metric space, and let $\mathbb{P}$ and $\mathbb{Q}$ represent Borel probability measures defined on $\mathcal{X}$. Given two sets of IID observations $S_{X}=\{\mathbf{x}^{(i)}\}_{i=1}^{n}$ and $S_{Z}=\{\mathbf{z}^{(i)}\}_{i=1}^{m}$ sampled from distributions $\mathbb{P}$ and $\mathbb{Q}$, respectively, kernel-based MMD (Borgwardt et al., 2006) quantifies the discrepancy between these two distributions:
$\textsc{MMD}(\mathbb{P},\mathbb{Q};\mathbb{H}_{k})$
$=\|\mu_{\mathbb{P}}-\mu_{\mathbb{Q}}\|_{\mathbb{H}_{k}}$ (2)
$=\sqrt{\mathbb{E}[k(X,X^{\prime})]+\mathbb{E}[k(Z,Z^{\prime})]-2\mathbb{E}[k(X,Z)]},$
where $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is the kernel of a reproducing kernel Hilbert space $\mathbb{H}_{k}$, $\mu_{\mathbb{P}}:=\mathbb{E}[k(\cdot,X)]$ and $\mu_{\mathbb{Q}}:=\mathbb{E}[k(\cdot,Z)]$ are the kernel mean embeddings of $\mathbb{P}$ and $\mathbb{Q}$, respectively.
For characteristic kernels, $\mu_{\mathbb{P}}=\mu_{\mathbb{Q}}$ implies $\mathbb{P}=\mathbb{Q}$, and thus, $\textsc{MMD}(\mathbb{P},\mathbb{Q};\mathcal{H}_{k})=0$ if and only if $\mathbb{P}=\mathbb{Q}$. In practice, we use the estimator from a recent work that can effectively measure the discrepancies between AEs and CEs (Gao et al., 2021), which is defined as:
$\widehat{\textsc{MMD}}^{2}_{\rm u}(S_{X},S_{Z};k_{\omega})=\frac{1}{n(n-1)}\sum_{i\neq j}H_{ij},$ (3)
where $H_{ij}=k_{\omega}(\mathbf{x}_{i},\mathbf{x}_{j})+k_{\omega}(\mathbf{z}_{i},\mathbf{z}_{j})-k_{\omega}(\mathbf{x}_{i},\mathbf{z}_{j})-k_{\omega}(\mathbf{z}_{i},\mathbf{x}_{j})$, and $k_{\omega}(\mathbf{x},\mathbf{z})$ is defined as:
$k_{\omega}(\mathbf{x},\mathbf{z})=\left[(1-\beta_{0})s_{\widehat{h_{\mathcal{C}}^{*}}}(\mathbf{x},\mathbf{z})+\beta_{0}\right]q(\mathbf{x},\mathbf{z}),$ (4)
where $\beta_{0}\in(0,1)$ and $q(\mathbf{x},\mathbf{z})$, i.e., the Gaussian kernel with bandwidth $\sigma_{q}$, are two important components ensuring that $k_{\omega}(\mathbf{x},\mathbf{z})$ serves as a characteristic kernel (Liu et al., 2020). Additionally, $s_{\widehat{h_{\mathcal{C}}^{*}}}(\mathbf{x},\mathbf{z})$ represents a deep kernel function designed to measure the similarity between $\mathbf{x}$ and $\mathbf{z}$ by utilizing semantic features extracted via the second last layer in $\widehat{h_{\mathcal{C}}^{*}}$ (i.e., a well-trained classifier on CEs). In practice, $s_{\widehat{h_{\mathcal{C}}^{*}}}(\mathbf{x},\mathbf{z})$ is a well-trained feature extractor (e.g., a classifier without the last layer).
4.2 Training Process of DDAD
In this section, we discuss the training process of DDAD, which includes optimizing MMD and training the denoiser. We provide detailed algorithmic descriptions for the training process of DDAD in Algorithms 1 and 2.
Optimizing MMD. Following Liu et al. (2020), the test power of MMD can be maximized by maximizing the following objective (i.e., optimize $k_{\omega}$):
$J(\mathbb{P},\mathbb{Q};k_{\omega})=\textsc{MMD}^{2}(\mathbb{P},\mathbb{Q};k_{\omega})/\sigma(\mathbb{P},\mathbb{Q};k_{\omega}),$
$\sigma(\mathbb{P},\mathbb{Q};k_{\omega}):=\sqrt{4(\mathbb{E}[H_{12}H_{13}]-\mathbb{E}[H_{12}]^{2})}$ and $H_{12},H_{13}$ refer to the $H_{ij}$ in Section 4.1. However, $J(\mathbb{P},\mathbb{Q};k_{\omega})$ cannot be directly optimized because $\textsc{MMD}^{2}(\mathbb{P},\mathbb{Q};k_{\omega})$ and $\sigma(\mathbb{P},\mathbb{Q};k_{\omega})$ depend on $\mathbb{P}$ and $\mathbb{Q}$ that are unknown. Therefore, instead, we can optimize an estimator of $J(\mathbb{P},\mathbb{Q};k_{\omega})$:
$\hat{J}_{\lambda}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega})$
$:=\widehat{\textsc{MMD}}^{2}_{\rm u}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega})/\hat{\sigma}_{\lambda}(S_{\mathcal{C}},S_{\mathcal{A}};k_{\omega}),$ (5)
where $S_{\mathcal{C}}$ are clean samples, $S_{\mathcal{A}}$ can be any adversarial samples, $\hat{\sigma}^{2}_{\lambda}$ is a regularized estimator of $\sigma^{2}$ and $\lambda$ is a small constant to avoid 0 division (here we assume $m=n$ to obtain the asymptotic distribution of the MMD estimator):
$\hat{\sigma}^{2}_{\lambda}$ $=\frac{4}{n^{3}}\sum^{n}_{i=1}\left(\sum^{n}_{j=1}H_{ij}\right)^{2}$
$-\frac{4}{n^{4}}\left(\sum^{n}_{i=1}\sum^{n}_{j=1}H_{ij}\right)^{2}+\lambda.$ (6)
We can obtain optimized $k_{\omega}$ (we denote it as $k^{*}_{\omega}$) by maximizing Eq. 4.2 on the training set. Then, we define MMD-OPT as the MMD estimator with an optimized characteristic kernel $k^{*}_{\omega}$:
[…]
where $S^{\prime}_{X}$ and $S^{\prime}_{Z}$ can be any two batches of samples from either the clean or the adversarial domain.
Training the denoiser. In this paper, we use DUNET (Liao et al., 2018) as our denosing model. To train the denoiser, we first randomly generate noise $\mathbf{n}$ from a Gaussian distribution $\mathbb{N}(\mu,\sigma^{2})$ and add $\mathbf{n}$ to $S_{\mathcal{A}}$ that are generated from clean data-label pairs $(S_{\mathcal{C}},Y_{\mathcal{C}})$, resulting in noise-injected AEs:
$S_{\text{noise}}=S_{\mathcal{A}}+\mathbf{n}.$
The design of injecting Gaussian noise is inspired by previous works showing that applying denoised smoothing to a denoiser-classifier pipeline can provide certified robustness (Salman et al., 2020b; Carlini et al., 2023). Following Lin et al. (2024), we set $\mu=0$ and $\sigma=0.25$ by default. Then, we can obtain denoised samples $S_{\text{denoised}}$ by feeding $S_{\text{noise}}$ to a denoiser $g$ with parameters ${\bm{\theta}}$:
$S_{\text{denoised}}=g_{\bm{\theta}}(S_{\text{noise}}).$
Ideally, $S_{\text{denoised}}$ should perform in the same way as its clean counterpart $S_{\mathcal{C}}$. To achieve this, motivated by our theoretical analysis in Section 3, the optimized parameters $\bm{\theta}^{*}$ are obtained by minimizing the distributional discrepancy towards the direction of making the classification correct, i.e., minimize MMD-OPT and the cross-entropy loss ${\mathcal{L}}_{\text{ce}}$ simultaneously:
$g_{\bm{\theta}^{*}}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\text{MMD-OPT}(S_{\mathcal{C}},g_{\bm{\theta}}(S_{\text{noise}}))$
$+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(g_{\bm{\theta}}(S_{\text{noise}})),Y_{\mathcal{C}}),$ (8)
where $\alpha>0$ is a regularization term ($10^{-2}$ by default) and $\widehat{h_{\mathcal{C}}^{*}}$ is the pre-trained classifier.
4.3 Inference Process of DDAD
In this section, we discuss the two-pronged inference process of DDAD in detail.
The use of validation data. In the inference phase, we define a batch of clean validation data as $S_{\mathcal{V}}$ and the test data as $S_{\mathcal{T}}$. In practice, $S_{\mathcal{V}}$ is extracted from the training data and is completely inaccessible during the training. Then $S_{\mathcal{V}}$ serves as a reference to measure the distributional discrepancy. According to Eq. 7, the distributional discrepancies between $S_{\mathcal{V}}$ and $S_{\mathcal{T}}$ can be defined as:
[…]
The two-pronged inference process. (1) if MMD-OPT($S_{\mathcal{V}}$, $S_{\mathcal{T}}$) in Eq. 9 is less than a threshold value $t$, i.e., MMD-OPT($S_{\mathcal{V}}$, $S_{\mathcal{T}}$) $<t$, then $S_{\mathcal{T}}$ will be treated as CEs and directly fed into the classifier. Then the output will be $\widehat{h^{*}_{\mathcal{C}}}(S_{\mathcal{T}})$, where $\widehat{h^{*}_{\mathcal{C}}}$ is a well-trained classifier; (2) otherwise, $S_{\mathcal{T}}$ will be treated as AEs and denoised by the denoiser. Then, the output will be $\widehat{h^{*}_{\mathcal{C}}}(g_{\bm{\theta}^{*}}(S_{\mathcal{T}}))$, where $g_{\bm{\theta}^{*}}$ is a well-trained denoiser by our method.
Table 1: Clean and robust accuracy (%) against adaptive white-box attacks (left: $\ell_{\infty}~{}(\epsilon=8/255)$, right: $\ell_{2}~{}(\epsilon=0.5)$) on CIFAR-10. means this method uses WideResNet-34-10. * means this method is trained with extra data. We report averaged results and standard deviations of our method for 5 runs. We show the most successful defense in bold.
$\ell_{\infty}~{}(\epsilon=8/255)$
Type Method Clean Robust
WRN-28-10
AT Gowal et al. (2021) 87.51 63.38
Gowal et al. (2020)* 88.54 62.76
Pang et al. (2022a) 88.62 61.04
AP Yoon et al. (2021) 85.66 33.48
Nie et al. (2022) 90.07 46.84
Lee & Kim (2023) 90.16 55.82
Ours DDAD 94.16 $\pm$ 0.08 67.53 $\pm$ 1.07
WRN-70-16
AT Rebuffi et al. (2021)* 92.22 66.56
Gowal et al. (2021) 88.75 66.10
Gowal et al. (2020)* 91.10 65.87
AP Yoon et al. (2021) 86.76 37.11
Nie et al. (2022) 90.43 51.13
Lee & Kim (2023) 90.53 56.88
Ours DDAD 93.91 $\pm$ 0.11 67.68 $\pm$ 0.87
$\ell_{2}~{}(\epsilon=0.5)$
Type Method Clean Robust
WRN-28-10
AT Rebuffi et al. (2021)* 91.79 78.80
Augustin et al. (2020) 93.96 78.79
Sehwag et al. (2022) 90.93 77.24
AP Yoon et al. (2021) 85.66 73.32
Nie et al. (2022) 91.41 79.45
Lee & Kim (2023) 90.16 83.59
Ours DDAD 94.16 $\pm$ 0.08 84.38 $\pm$ 0.81
WRN-70-16
AT Rebuffi et al. (2021)* 95.74 82.32
Gowal et al. (2020)* 94.74 80.53
Rebuffi et al. (2021) 92.41 80.42
AP Yoon et al. (2021) 86.76 75.66
Nie et al. (2022) 92.15 82.97
Lee & Kim (2023) 90.53 83.57
Ours DDAD 93.91 $\pm$ 0.11 84.03 $\pm$ 0.75
Table 2: Clean and robust accuracy (%) against adaptive white-box attacks $\ell_{\infty}~{}(\epsilon=4/255)$ on ImageNet-1K. We report averaged results and standard deviations of our method for 3 runs. We show the most successful defense in bold.
$\ell_{\infty}~{}(\epsilon=4/255)$
Type Method Clean Robust
RN-50
AT Salman et al. (2020a) 64.02 34.96
Engstrom et al. (2019) 62.56 29.22
Wong et al. (2020) 55.62 26.24
AP Nie et al. (2022) 71.48 38.71
Lee & Kim (2023) 70.74 42.15
Ours DDAD 78.61 $\pm$ 0.04 53.85 $\pm$ 0.23
5 Experiments
5.1 Experiment Settings
We briefly introduce the experiment settings here and provide a more detailed version in Appendix D.
Dataset and target models. We evaluate DDAD on two benchmark datasets with different scales, i.e., CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009). For the target models, we mainly use three architectures with different capacities: ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016) and Swin Transformer (Liu et al., 2021).
Baseline settings. DDAD is a two-pronged adversarial defense method, which is different from most existing defense methods. In terms of the pipeline structure, MagNet (Meng & Chen, 2017) is the only similar defense method to ours, which also contains a two-pronged process. However, MagNet is now considered outdated, making it unfair for DDAD to compare with it. Therefore, to make the comparison as fair as possible, we follow a recent study on robust evaluation (Lee & Kim, 2023) to compare our method with SOTA adversarial training (AT) methods in RobustBench (Croce et al., 2020) and adversarial purification (AP) methods selected by Lee & Kim (2023).
Evaluation settings. We mainly use PGD+EOT (Athalye et al., 2018b) and AutoAttack (Croce & Hein, 2020a) to compare our method with different baseline methods. Specifically, following Lee & Kim (2023), we evaluate AP methods on the PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We set the EOT iteration to 20 for both datasets. We evaluate AT baseline methods using AutoAttack with 100 update iterations, as AT methods have seen PGD attacks during training, leading to overestimated results when evaluated on PGD+EOT (Lee & Kim, 2023). For our method, we implicitly design an adaptive white-box attack by considering the entire defense mechanism of DDAD (see Algorithm 3). To make a fair comparison, we evaluate our method on both adaptive white-box PGD+EOT attack and adaptive white-box AutoAttack with the same configurations mentioned above. Notably, we find that our method achieves the worst-case robust accuracy on adaptive white-box PGD+EOT attack. Therefore, we report the robust accuracy of our method on adaptive white-box PGD+EOT attack for Table 1 and 2. On CIFAR-10, the maximum allowed perturbaiton budget $\epsilon$ for $\ell_{\infty}$-norm-based attacks and $\ell_{2}$-norm-based attacks is set to $8/255$ and $0.5$, respectively. While on ImageNet-1K, we set $\epsilon=4/255$ for $\ell_{\infty}$-norm-based attacks.
Implementation details of DDAD. To avoid the evaluation bias caused by seeing similar attacks beforehand during training, we train both the MMD-OPT and the denoiser using $\ell_{\infty}$-norm MMA attack (Gao et al., 2022), which differs significantly from PGD+EOT and AutoAttack. Then, we use unseen attacks to evaluate DDAD. For optimizing the MMD, following Gao et al. (2021), we set the learning rate to be $2\times 10^{-4}$ and the epoch number to be 200. For training the denoiser, we set the epoch number to be 60. The initial learning rate is set to $1\times 10^{-3}$ for both datasets and is divided by 10 at the 45th and 60th epoch to avoid robust overfitting (Rice et al., 2020). We provide more detailed implementation details in Appendix D.
5.2 Defending against Adaptive White-box Attacks
Result analysis on CIFAR-10. Table 1 shows the evaluation performance of DDAD against adaptive white-box PGD+EOT attack with $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ on CIFAR-10. Compared to SOTA defense methods, DDAD improves clean and robust accuracy by a notable margin. The evaluation results against BPDA+EOT on CIFAR-10 can be found in Appendix E.1.
Result analysis on ImageNet-1K. Table 2 shows the evaluation performance of DDAD against adaptive white-box PGD+EOT attack with $\ell_{\infty}(\epsilon=4/255)$ on ImageNet-1K. The advantages of our method over baselines become more significant on large-scale datasets. Specifically, compared with AP methods that rely on density estimation (Nie et al., 2022; Lee & Kim, 2023), our method improves clean accuracy by at least 7.13% and robust accuracy by 11.70% on ResNet-50. This empirical evidence supports that identifying distributional discrepancies is a simpler and more feasible task than estimating data density, especially on large-scale datasets such as ImageNet-1K.
5.3 Defending against Unseen Transfer Attacks
Since DDAD requires AEs to train the MMD-OPT and the denoiser, it is important for us to evaluate the transferability of our method. We report the transferability of our method (trained on WideResNet-28-10) under different threat models, which include WideResNet-70-16, ResNet-18, ResNet-50 and Swin Transformer in Appendix E.2. We use PGD+EOT $\ell_{\infty}$ and C&W $\ell_{2}$ (Carlini & Wagner, 2017) with 200 iterations for evaluation. Experiment results show that our method generalizes well to unseen transfer attacks.
Figure 2: Top: clean accuracy (%) vs. batch size; Bottom: mixed accuracy (%) vs. proportion of AEs in a batch (%). We plot averaged results and standard deviations of 5 runs.
5.4 Ablation Studies
Ablation study on batch size. Identifying distributional discrepancies requires the data to be processed in batches. Therefore, we aim to determine how much data in a batch will not affect the stability of our method. Figure 2 (top) shows the clean accuracy of our method on CIFAR-10 with different batch sizes, ranging from 10 to 110. We find that once the batch size exceeds 100, the performance of our method is stable. Therefore, in this paper, we set the test batch size to 100 for all evaluations.
Ablation study on mixed data batches. We explore a more challenging scenario for our method, in which each data batch contains a mixture of CEs and AEs. Figure 2 (bottom) shows the mixed accuracy (i.e., the accuracy on mixed data) of our method on CIFAR-10 with different proportions of AEs (generated by adaptive white-box PGD+EOT $\ell_{\infty}$ with $\epsilon=8/255$) in each batch, ranging from 0% (i.e., pure CEs) to 100% (i.e., pure AEs). Initially, the mixed accuracy drops from over 90% to approximately 80%. This is because, with a high proportion of CEs, the MMD-OPT has a high chance to regard the entire batch as clean data. After that (i.e., from 30% onwards), the mixed accuracy degrades gradually to approximately 70%. This is because, as the proportion of AEs increases, the MMD-OPT regards the entire batch as adversarial and feeds it into the denoiser. Notably, DDAD can still outperform baseline methods (see Appendix E.3).
Ablation study on threshold values of MMD-OPT. We explore the impact of threshold values of MMD-OPT in Appendix E.4. We select the threshold based on the experimental results on the validation data. Specifically, a threshold value of $0.5$ is selected for CIFAR-10 and $0.02$ is selected for ImageNet-1K. It is reasonable to use a smaller threshold for ImageNet-1K because the distribution of AEs with $\epsilon=4/255$ (i.e., AEs for ImageNet-1K) will be closer to CEs than AEs with $\epsilon=8/255$ (i.e., AEs for CIFAR-10). Intuitively, when $\epsilon$ decreases to 0, AEs are the same as CEs (i.e., the distribution of AEs and CEs will be the same).
Ablation study on injecting Gaussian noise. We provide evaluation results of our method against adaptive white-box PGD+EOT attack with and without injecting Gaussian noise on CIFAR-10 in Appendix E.5. We find that injecting Gaussian noise can make DDAD generalize better.
Ablation study on the two-pronged process. We provide evaluation results of our method against adaptive white-box PGD+EOT attack with and without MMD-OPT on CIFAR-10 in Appendix E.6. We find that using the two-pronged process can largely improve clean accuracy.
5.5 Combination with Adversarial Training
We also combine several well-known AT methods (Madry et al., 2018; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020) with DDAD: detected CEs are directly classified by the AT-based classifier, while detected AEs are first denoised by our method before being classified. Experimental results show that combining AT methods with DDAD further enhances adversarial robustness (see Appendix E.7).
5.6 Compute Resource of DDAD
We report the compute resources used for training and evaluating DDAD in Appendix E.8. Compared to AT baselines, DDAD offers better training efficiency (e.g., it can scale to large datasets like ImageNet-1K). Additionally, although DDAD requires training an extra denoiser and MMD-OPT, it significantly outperforms AP baselines in inference speed. Furthermore, relying on a pre-trained generative model is not always feasible, as training such models at scale can be highly resource-intensive. Therefore, in general, DDAD provides a more lightweight design.
6 Related Work
We briefly review the related work here, and a more detailed version can be found in Appendix F.
Statistical adversarial data detection. Recently, statistical adversarial data detection (SADD) has attracted increasing attention in defending against AEs. For example, Gao et al. (2021) demonstrate that maximum mean discrepancy (MMD) is aware of adversarial attacks and leverage the distributional discrepancy between AEs and CEs to filter out AEs, which has been shown effective against unseen attacks. Based on this, Zhang et al. (2023) further propose expected perturbation score to measure the expected score of a sample after multiple perturbations.
Denoiser-based adversarial defense. Denoiser-based adversarial defense often leverages generative models to shift AEs back to their clean counterparts before feeding them into a classifier. In most literature, it is called adversarial purification (AP). At the early stage of AP, Meng & Chen (2017) propose a two-pronged defense called MagNet to remove adversarial noise by first using a detector to discard the detected AEs, and then using an autoencoder to purify the remaining samples. The following studies mainly focus on exploring the use of more powerful generative models for AP (Liao et al., 2018; Samangouei et al., 2018; Song et al., 2018; Yoon et al., 2021; Nie et al., 2022). Recently, the outstanding denoising capabilities of pre-trained diffusion models have been leveraged to purify AEs (Nie et al., 2022; Lee & Kim, 2023). The success of recent AP methods often relies on the assumption that there will be a pre-trained generative model that can precisely estimate the probability density of the CEs (Nie et al., 2022; Lee & Kim, 2023). However, even powerful generative models (e.g., diffusion models) may have an inaccurate density estimation, leading to unsatisfactory performance (Chen et al., 2024). By contrast, instead of estimating probability densities, our method directly minimizes the distributional discrepancies between AEs and CEs, leveraging the fact that identifying distributional discrepancies is simpler and more feasible.
7 Discussions on Batch-wise Detections
We briefly discuss the limitation, solutions and practicability here, and see Appendix G for detailed discussions.
Limitation and its solutions for user inference. DDAD leverages statistics based on distributional discrepancies, which requires the data to be processed in batches for adversarial detection. When the batch size is too small, the stability of DDAD will be affected (see Figure 2). To address this issue, each single sample provided by the user can be dynamically stored in a queue. Once the queue accumulates enough samples to form a batch, our method can then process the batch collectively using the proposed approach. A direct cost of this solution is the waiting time, as the system must accumulate enough samples (e.g., 50 samples) to form a batch before processing. However, in scenarios where data arrives quickly, the waiting time is typically very short, making this approach feasible for many real-time applications. Overall, it is a trade-off problem: using our method for user inference can obtain high robustness, but the cost is to wait for batch processing. Based on the performance improvements our method obtains over the baseline methods, we believe the cost is acceptable. Another possible solution is to find more robust statistics that can measure distributional discrepancies with fewer samples. We leave finding such statistics as future work.
Practicability beyond user inference. Other than user inference, our method is suitable for cleaning the data before fine-tuning the underlying model. In many domains, obtaining large quantities of high-quality data is challenging due to factors such as cost, privacy concerns, or the rarity of specific data. As a result, all possible samples with clean information are critical in these data-scarce domains. Then, a practical scenario is that there exists a pre-trained model on a large-scale dataset (e.g., a DNN trained on ImageNet-1K) and clients want to fine-tune the model to perform well on downstream tasks. If the data for downstream tasks contain AEs, our method can be applied to batch-wisely clean the data before fine-tuning the underlying model.
8 Conclusion
SADD-based defense methods empirically show that leveraging the distributional discrepancies can effectively defend against adversarial attacks. However, a potential limitation of SADD-based methods is that they will discard data batches that contain AEs, leading to the loss of clean information. To solve this problem, inspired by our theoretical analysis that minimizing distributional discrepancy can help reduce the expected loss on AEs, we propose a two-pronged adversarial defense called Distributional-Discrepancy-based Adversarial Defense (DDAD) that leverages the effectiveness of SADD-based methods without discarding input data. Extensive experiments demonstrate the effectiveness of DDAD against various adversarial attacks. In general, we hope this simple yet effective method could open up a new perspective on adversarial defenses.
Impact Statement
This study on adversarial defense mechanisms raises important ethical considerations that we have carefully addressed. We have taken steps to ensure our adversarial defense method is fair. We use widely accepted public benchmark datasets to ensure comparability of our results. Our evaluation encompasses a wide range of attack types and strengths to provide a comprehensive assessment of our defense mechanism. We have also carefully considered the broader impacts of our work. The proposed defense algorithm contributes to the development of more robust machine learning models, potentially improving the reliability of AI systems in various applications. We will actively engage with the research community to promote responsible development and use of adversarial defenses.
Appendix A Proof of Theorem 1
See 1
Proof.
Let $\phi_{\mathcal{C}}$ and $\phi_{A}$ be the density functions of $\mathcal{D_{C}}$ and $\mathcal{D_{A}}$:
$R(h,f_{\mathcal{A}},\mathcal{D_{A}})$ $=R(h,f_{\mathcal{A}},\mathcal{D_{A}})+R(h,f_{\mathcal{C}},\mathcal{D_{C}})-R(h,f_{\mathcal{C}},\mathcal{D_{C}})+R(h,f_{\mathcal{A}},\mathcal{D_{C}})-R(h,f_{\mathcal{A}},\mathcal{D_{C}})$
$\leq R(h,f_{\mathcal{C}},\mathcal{D_{C}})+|R(h,f_{\mathcal{A}},\mathcal{D_{C}})-R(h,f_{\mathcal{C}},\mathcal{D_{C}})|+|R(h,f_{\mathcal{A}},\mathcal{D_{A}})-R(h,f_{\mathcal{A}},\mathcal{D_{C}})|$
$\leq R(h,f_{\mathcal{C}},\mathcal{D_{C}})+\mathbb{E}\left[|f_{\mathcal{C}}(\mathbf{x})-f_{\mathcal{A}}(\mathbf{x})|\right]+|R(h,f_{\mathcal{A}},\mathcal{D_{A}})-R(h,f_{\mathcal{A}},\mathcal{D_{C}})|$
[…]
$\overset{(a)}{\leq}R(h,f_{\mathcal{C}},\mathcal{D_{C}})+\mathbb{E}\left[|f_{\mathcal{C}}(\mathbf{x})-f_{\mathcal{A}}(\mathbf{x})|\right]+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})$
=$\overset{(b)}{=}R(h,f_{\mathcal{C}},\mathcal{D_{C}})+\mathbb{E}\left[|f_{\mathcal{C}}(\mathbf{x})-f_{\mathcal{C}}(\mathbf{x})|\right]+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}})$
$=R(h,f_{\mathcal{C}},\mathcal{D_{C}})+d_{1}(\mathcal{D_{C}},\mathcal{D_{A}}),$
where $(a)$ is based on Definition 1 and $(b)$ is based on Corollary 1. ∎
Appendix B Mathematical Notations in Section 4
$\mathcal{X}$ A separable metric space in $\mathbb{R}^{d}$
$\mathbb{P},\mathbb{Q}$ Borel probability measures defined on $\mathcal{X}$
$S_{X}$ $n$ IID observations sampled from $\mathbb{P}$, i.e., $\{\mathbf{x}^{(i)}\}_{i=1}^{n}$
$S_{Z}$ $m$ IID observations sampled from $\mathbb{Q}$, i.e., $\{\mathbf{z}^{(i)}\}_{i=1}^{m}$
$\mathbb{H}_{k}$ A reproducing kernel Hilbert space
$k_{\omega}$ A kernel of $\mathbb{H}_{k}$ with parameters $\omega$
$\mu_{\mathbb{P}}$ The kernel mean embedding of $\mathbb{P}$
$\mu_{\mathbb{Q}}$ The kernel mean embedding of $\mathbb{Q}$
$H_{ij}$ $k_{\omega}(\mathbf{x}_{i},\mathbf{x}_{j})+k_{\omega}(\mathbf{z}_{i},\mathbf{z}_{j})-k_{\omega}(\mathbf{x}_{i},\mathbf{z}_{j})-k_{\omega}(\mathbf{z}_{i},\mathbf{x}_{j})$
$s_{\widehat{h^{*}_{\mathcal{C}}}}$ A deep kernel function that measures the similarity between $\mathbf{x}$ and $\mathbf{z}$
$\widehat{h^{*}_{\mathcal{C}}}$ A well-trained classifier
$\beta_{0}$ A constant $\in(0,1)$
$q$ The Gaussian kernel with bandwidth $\sigma_{q}$
$J$ The objective function of optimizing MMD
$\mu$, $\sigma$ Mean and standard deviation
$\lambda$ A small constant to avoid 0 division
$\mathbf{n}$ Gaussian noise, i.e., $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$
$g_{\bm{\theta}}$ A denoiser with parameters $\bm{\theta}$
$S_{\mathcal{C}}$ Clean samples
$Y_{\mathcal{C}}$ Ground truth labels of $S_{\mathcal{C}}$
$S_{\mathcal{A}}$ Adversarial examples
$S_{\mathcal{\text{noise}}}$ Noise-injected adversarial examples
$S_{\mathcal{\text{denoised}}}$ Denoised samples
$\alpha$ A regularization term
Appendix C Algorithms
1:  Input: clean data $S^{\text{train}}_{\mathcal{C}}$, adversarial data $S^{\text{train}}_{\mathcal{A}}$, learning rate $\eta$, epoch $T$;
2:  Initialize $\omega\leftarrow\omega_{0}$; $\lambda\leftarrow 10^{-8}$;
3:  for epoch $=1,...,T$ do
4:     $S^{\prime}_{\mathcal{C}}\leftarrow$ minibatch from $S^{\text{train}}_{\mathcal{C}}$;
5:     $S^{\prime}_{\mathcal{A}}\leftarrow$ minibatch from $S^{\text{train}}_{\mathcal{A}}$;
6:     $k_{\omega}\leftarrow$ kernel function with parameters $\omega$ using Eq. 4;
7:     $M(\omega)\leftarrow\widehat{\textsc{MMD}}^{2}_{\rm u}(S^{\prime}_{\mathcal{C}},S^{\prime}_{\mathcal{A}};k_{\omega})$ using Eq. 3;
8:     $V_{\lambda}(\omega)\leftarrow\hat{\sigma}_{\lambda}(S^{\prime}_{\mathcal{C}},S^{\prime}_{\mathcal{A}};k_{\omega})$ using Eq. 4.2;
9:     $\hat{J}_{\lambda}(\omega)\leftarrow M(\omega)/\sqrt{V_{\lambda}(\omega)}$ using Eq. 4.2;
10:     $\omega\leftarrow\omega+\eta\nabla_{\text{Adam}}\hat{J}_{\lambda}(\omega)$;
11:  end for
12:  Output: $k^{*}_{\omega}$
Algorithm 1 Optimizing MMD (Liu et al., 2020).
1:  Input: clean data-label pairs $(S^{\text{train}}_{\mathcal{C}},Y^{\text{train}}_{\mathcal{C}})$, optimized characteristic kernel $k^{*}_{\omega}$ by Algorithm 1, pre-trained classifier $\widehat{h_{\mathcal{C}}^{*}}$, denoiser $g$ with parameters $\bm{\theta}$, learning rate $\eta$, epoch $T$;
2:  Initialize $\mu\leftarrow 0$; $\sigma\leftarrow 0.25$; $\alpha\leftarrow 10^{-2}$;
3:  for epoch $=1,...,T$ do
4:     $(S^{\prime}_{\mathcal{C}},Y^{\prime}_{\mathcal{C}})\leftarrow$ minibatch from $(S^{\text{train}}_{\mathcal{C}},Y^{\text{train}}_{\mathcal{C}})$;
5:     $S^{\prime}_{\mathcal{A}}\leftarrow$ adversarial examples generated from $(S^{\prime}_{\mathcal{C}},Y^{\prime}_{\mathcal{C}})$;
6:     generate Gaussian noise: $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$;
7:     $S^{\prime}_{\text{noise}}\leftarrow S^{\prime}_{\mathcal{A}}+\mathbf{n}$;
[…]
9:     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\text{Adam}}(\text{MMD-OPT}(S^{\prime}_{\mathcal{C}},g_{\bm{\theta}}(S^{\prime}_{\text{noise}}))+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(g_{\bm{\theta}}(S^{\prime}_{\text{noise}})),Y^{\prime}_{\mathcal{C}}))$ using Eq. 4.2;
10:  end for
11:  Output: denoiser $g$ with well-trained parameters $\bm{\theta}^{*}$
Algorithm 2 Training the denoiser.
Algorithm 3 Adaptive white-box PGD+EOT attack for DDAD.
1:  Input: clean data-label pairs $(S_{\mathcal{C}},Y_{\mathcal{C}})$, optimized characteristic kernel $k^{*}_{\omega}$ by Algorithm 1, pre-trained classifier $\widehat{h_{\mathcal{C}}^{*}}$, denoiser $g$ with parameters $\bm{\theta}$, maximum allowed perturbation $\epsilon$, step size $\eta$, PGD iteration $T$, EOT iteration $K$;
2:  Initialize adversarial data $S_{\mathcal{A}}\leftarrow S_{\mathcal{C}}$; $\mu\leftarrow 0$; $\sigma\leftarrow 0.25$; $\alpha\leftarrow 10^{-2}$; $t\leftarrow 0.05$;
3:  for PGD iteration $1,...,T$ do
4:     Initialize gradients over EOT ${\mathcal{G}}_{\text{EOT}}\leftarrow\mathbf{0};$
[…]
6:     for EOT iteration $1,...,K$ do
[…]
8:           ${\mathcal{G}}_{\text{EOT}}\leftarrow{\mathcal{G}}_{\text{EOT}}+\nabla_{S_{\mathcal{A}}}(\text{MMD-OPT}(S_{\mathcal{C}},S_{\mathcal{A}})+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(S_{\mathcal{A}}),Y_{\mathcal{C}}))$;
9:        else
10:           Generate Gaussian noise: $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$; $S_{\text{noise}}\leftarrow S_{\mathcal{A}}+\mathbf{n}$;
11:           ${\mathcal{G}}_{\text{EOT}}\leftarrow{\mathcal{G}}_{\text{EOT}}+\nabla_{S_{\mathcal{A}}}(\text{MMD-OPT}(S_{\mathcal{C}},S_{\mathcal{A}})+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(g_{\bm{\theta}}(S_{\text{noise}})),Y_{\mathcal{C}}))$;
12:        end if
13:     end for
14:     ${\mathcal{G}}_{\text{EOT}}\leftarrow\frac{1}{K}{\mathcal{G}}_{\text{EOT}}$;
15:     Update adversarial data $S_{\mathcal{A}}\leftarrow\Pi_{\mathcal{B}_{\epsilon}(S_{\mathcal{C}})}\left(S_{\mathcal{A}}+\eta\cdot\text{sign}({\mathcal{G}}_{\text{EOT}})\right)$;
16:  end for
17:  Output: $S_{\mathcal{A}}$
Appendix D Detailed Experiment Settings
D.1 Dataset and target models
We evaluate the effectiveness of DDAD on two benchmark datasets with different scales, i.e., CIFAR-10 (Krizhevsky et al., 2009) (small scale) and ImageNet-1K (Deng et al., 2009) (large scale). Specifically, CIFAR-10 contains 50,000 training images and 10,000 test images, divided into 10 classes. ImageNet-1K is a large-scale dataset that contains 1,000 classes and consists of 1,281,167 training images, 50,000 validation images, and 100,000 test images. For the target models, we use three widely used architectures with different scales: ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016) and Swin Transformer (Liu et al., 2021). Specifically, following Lee & Kim (2023), we use WideResNet-28-10 and WideResNet-70-16 to evaluate the performance of defense methods on CIFAR-10 and we use ResNet-50 to evaluate the performance of defense methods on ImageNet-1K. Additionally, we examine the transferability of our method under different threat models, which include ResNet-18, ResNet-50, WideResNet-70-16 and Swin Transformer.
D.2 Baseline Settings
DDAD is a two-pronged adversarial defense method, which is different from most existing defense methods. In terms of the pipeline structure, MagNet (Meng & Chen, 2017) is the only similar defense method to ours, which also contains a two-pronged process. However, MagNet is now considered outdated, making it unfair for DDAD to compare with it. Therefore, to make the comparison as fair as possible, we follow a recent study on robust evaluation (Lee & Kim, 2023) to compare our method with SOTA adversarial training (AT) methods in RobustBench (Croce et al., 2020) and adversarial purification (AP) methods selected by Lee & Kim (2023).
D.3 Evaluation Settings
We mainly use PGD+EOT (Athalye et al., 2018b) and AutoAttack (Croce & Hein, 2020a) to compare our method with different baseline methods. Specifically, following Lee & Kim (2023), we evaluate AP methods on the PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We set the EOT iteration to 20 for both datasets. We evaluate AT baseline methods using AutoAttack with 100 update iterations, as AT methods have seen PGD attacks during training, leading to overestimated results when evaluated on PGD+EOT (Lee & Kim, 2023). For our method, we implicitly design an adaptive white-box attack by considering the entire defense mechanism of DDAD. To make a fair comparison, we evaluate our method on both adaptive white-box PGD+EOT attack and adaptive white-box AutoAttack with the same configurations mentioned above. Notably, we find that our method achieves the worst-case robust accuracy on adaptive white-box PGD+EOT attack. Therefore, we report the robust accuracy of our method on adaptive white-box PGD+EOT attack for Table 1 and 2. The algorithmic descriptions of the adaptive white-box attack is provided in Algorithm 3. On CIFAR-10, $\epsilon$ for $\ell_{\infty}$-norm-based attacks and $\ell_{2}$-norm-based attacks is set to $8/255$ and $0.5$, respectively. While on ImageNet-1K, we set $\epsilon=4/255$ for $\ell_{\infty}$-norm-based attacks. We also evaluate our method against BPDA+EOT (Hill et al., 2021) on CIFAR-10. For BPDA+EOT, we use the implementation of Hill et al. (2021) with default hyperparameters for evaluation. For transferability experiments, we use PGD+EOT $\ell_{\infty}$ and C&W $\ell_{2}$ (Carlini & Wagner, 2017) for evaluation. Specifically, the iteration number of C&W $\ell_{2}$ is set to 200. For $\ell_{\infty}$-norm transfer attacks, we examine the robustness of our method under $\epsilon=8/255$ and $\epsilon=12/255$. For C&W $\ell_{2}$, we examine our method under $\epsilon=0.5$ and $\epsilon=1.0$.
D.4 Implementation Details of DDAD
To avoid the evaluation bias caused by learning similar attacks beforehand during training, we train both the MMD-OPT and the denoiser using the MMA attack with $\ell_{\infty}$-norm (Gao et al., 2022), which differs significantly from PGD+EOT and AutoAttack. Then, we use unseen attacks to evaluate DDAD. We set $\epsilon=8/255$ with a step size of $2/255$ for CIFAR-10, and $\epsilon=4/255$ with a step size of $1/255$ for ImageNet-1K. For optimizing the MMD, following Gao et al. (2021), we set the learning rate to be $2\times 10^{-4}$ and the epoch number to be 200. For training the denoiser, we set the initial learning rate to $1\times 10^{-3}$ for both CIFAR-10 and ImageNet-1K. We set the epoch number to be 60 and divide the learning rate by 10 at the 45th epoch and 60th epoch to avoid robust overfitting (Rice et al., 2020). The training batch size is set to 500 for CIFAR-10 and 128 for ImageNet-1K. The optimizer we use is Adam (Kingma & Ba, 2015). To improve the training efficiency on ImageNet-1K, we randomly select 100 samples from each class, resulting in 100,000 training samples in total. Notably, during the inference time, we evaluate our method using the entire testing set for both CIFAR-10 and ImageNet-1K. The batch size for evaluation is set to 100 for all datasets.
Appendix E Additional Experiments
E.1 Defending against BPDA+EOT Attack
Table 3: Clean (%) and robust accuracy (%) of defense methods against BPDA+EOT attack under $\ell_{\infty}(\epsilon=8/255)$ threat models on CIFAR-10. We report averaged results and standard deviations of DDAD for 5 runs. We show the most successful defense in bold.
Category Model Method Clean Robust Average
Adversarial Training RN-18 Madry et al. (2018) 87.30 45.80 66.55
Zhang et al. (2019) 84.90 45.80 65.35
WRN-28-10 Carmon et al. (2019) 89.67 63.10 76.39
Gowal et al. (2020) 89.48 64.08 77.28
Adversarial Purification RN-18 Yang et al. (2019) 94.80 40.80 67.80
RN-62 Song et al. (2018) 95.00 9.00 52.00
WRN-28-10 Hill et al. (2021) 84.12 54.90 69.51
Yoon et al. (2021) 86.14 70.01 78.08
Wang et al. (2022) 93.50 79.83 86.67
Nie et al. (2022) 89.02 81.40 85.21
Lee & Kim (2023) 90.16 88.40 89.28
Ours WRN-28-10 DDAD 94.16 $\pm$ 0.08 87.13 $\pm$ 1.19 90.65
E.2 Defending against Unseen Transfer Attacks
Table 4: Robust accuracy (%) of DDAD trained on WideResNet-28-10 against unseen transfer attacks on CIFAR-10. Attackers cannot access parameters of WideResNet-28-10, thereby it is in a gray-box setting. We report averaged results and standard deviations of 5 runs.
Trained on WRN-28-10
Unseen Transfer Attack WRN-70-16 RN-18 RN-50 Swin-T
PGD+EOT ($\ell_{\infty}$) $\epsilon=8/255$ 80.84 $\pm$ 0.46 80.78 $\pm$ 0.60 81.47 $\pm$ 0.30 81.46 $\pm$ 0.29
$\epsilon=12/255$ 80.26 $\pm$ 0.60 80.54 $\pm$ 0.45 80.98 $\pm$ 0.36 80.40 $\pm$ 0.41
C&W ($\ell_{2}$) $\epsilon=0.5$ 82.45 $\pm$ 0.19 91.30 $\pm$ 0.20 89.26 $\pm$ 0.11 93.45 $\pm$ 0.17
$\epsilon=1.0$ 81.20 $\pm$ 0.39 90.37 $\pm$ 0.17 88.65 $\pm$ 0.22 93.41 $\pm$ 0.18
E.3 Ablation Study on Mixed Data Batches
Table 5: Mixed accuracy (%) of defense methods against adaptive white-box attacks $\ell_{\infty}(\epsilon=8/255)$ on CIFAR-10 under different proportions of AEs. The target model is WRN-28-10. We report averaged results and standard deviations of 5 runs. We show the most successful defense in bold.
Method Proportion of AEs in Each Batch (%)
10 20 30 40 50 60 70 80 90 100
Rebuffi et al. (2021) 85.10 82.68 80.27 77.86 75.45 73.03 70.62 68.21 65.79 63.38
Augustin et al. (2020) 85.96 83.38 80.81 78.23 75.65 73.07 70.49 67.92 65.34 62.76
Sehwag et al. (2022) 85.86 83.10 80.35 77.59 74.83 72.07 69.31 66.56 63.80 61.04
Yoon et al. (2021) 81.80 76.83 71.87 66.90 61.94 56.97 52.01 47.04 42.08 37.11
Nie et al. (2022) 85.75 81.42 77.10 72.78 68.46 64.13 59.81 55.49 55.16 46.84
Lee & Kim (2023) 86.73 83.29 79.86 76.42 72.99 69.56 66.12 62.69 59.25 55.82
91.22 87.15 81.77 79.94 77.78 76.14 74.22 72.37 69.56 67.53
Ours $\pm$ 0.47 $\pm$ 0.58 $\pm$ 0.66 $\pm$ 0.66 $\pm$ 0.51 $\pm$ 0.69 $\pm$ 0.53 $\pm$ 0.74 $\pm$ 0.83 $\pm$ 1.07
E.4 Ablation Study on Threshold Values of MMD-OPT
Table 6: Sensitivity of DDAD to the threshold values of MMD-OPT on CIFAR-10. We report clean and robust accuracy (%) against adaptive white-box attacks ($\epsilon=8/255$). The classifier used is WRN-28-10.
Threshold Value Clean PGD+EOT AutoAttack
$\ell_{\infty}$ $\ell_{2}$ $\ell_{\infty}$ $\ell_{2}$
0.05 94.16 66.98 73.40 72.21 85.96
0.07 94.16 66.98 73.40 72.21 85.96
0.10 94.16 66.98 73.40 72.21 85.96
0.50 94.16 66.98 84.38 72.21 85.96
0.70 94.16 66.98 84.38 72.21 85.96
1.00 94.16 64.75 84.38 72.21 85.96
Table 7: Sensitivity of DDAD to the threshold values of MMD-OPT on ImageNet-1K. We report clean and robust accuracy (%) against adaptive white-box attacks ($\epsilon=4/255$). The classifier used is RN-50.
Threshold Value Clean $\ell_{\infty}$)
0.010 76.61 53.75
0.015 76.61 53.75
0.020 78.61 53.75
0.025 78.61 53.75
0.030 78.61 0.46
0.040 78.61 0.46
0.050 78.61 0.46
E.5 Ablation Study on Injecting Gaussian Noise
Table 8: Robust accuracy (%) of our method with and without injecting Guassian noise against adaptive white-box PGD+EOT $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ on CIFAR-10. We report averaged results and standard deviations of 5 runs. We show the most successful defense in bold.
Gaussian Noise Model PGD+EOT $(\ell_{\infty})$ PGD+EOT $(\ell_{2})$
WRN-28-10 65.31 $\pm$ 0.67 81.04 $\pm$ 0.52
67.53 $\pm$ 1.07 84.38 $\pm$ 0.81
E.6 Ablation Study on the Two-pronged Process
Table 9: Clean and robust accuracy (%) of our method with and without the two-pronged process against adaptive white-box PGD+EOT $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ on CIFAR-10. We report averaged results and standard deviations of 5 runs. We show the most successful defense in bold.
Module Model Clean PGD+EOT $(\ell_{\infty})$ PGD+EOT $(\ell_{2})$
Denoiser only WRN-28-10 85.07 $\pm$ 0.16 71.76 $\pm$ 0.65 85.01 $\pm$ 0.50
Denoiser + MMD-OPT 94.16 $\pm$ 0.08 67.53 $\pm$ 1.07 84.37 $\pm$ 0.81
E.7 Combination with Adversarial Training
Table 10: Mixed accuracy (%) our method with and without different adversarial training methods under different proportions of AEs in each batch and different batch sizes against PGD+EOT $\ell_{\infty}(\epsilon=8/255)$. We show the most successful defense in bold.
Method Proportion of AEs in Each Batch (%)
10 20 30 40 50 60 70 80 90 100
Batch Size = 25
Vanilla AT (Madry et al., 2018) 71.89 68.80 65.72 62.63 59.55 56.46 53.38 50.29 47.21 44.12
+ Ours 72.81 69.08 66.32 62.67 60.69 56.98 54.60 50.82 48.87 45.75
TRADES (Zhang et al., 2019) 70.29 67.70 65.11 62.52 59.92 57.33 54.74 52.15 49.56 46.97
+ Ours 70.99 67.73 65.67 63.46 60.88 57.87 55.80 52.92 50.98 48.10
MART (Wang et al., 2020) 68.63 66.25 63.86 61.47 59.08 56.70 54.31 51.92 49.54 47.15
+ Ours 69.34 66.67 64.34 61.78 60.05 57.27 55.16 52.48 51.10 48.46
TRADES-AWP (Wu et al., 2020) 69.52 67.22 64.92 62.62 60.32 58.02 55.72 53.42 51.12 48.82
+ Ours 70.22 66.98 65.56 62.46 61.12 58.18 57.03 54.65 53.40 49.99
Batch Size = 50
Vanilla AT (Madry et al., 2018) 71.89 68.81 65.73 62.65 59.57 56.49 53.41 50.33 47.25 44.17
+ Ours 72.84 69.57 66.55 62.93 60.74 57.12 53.88 51.21 48.82 46.91
TRADES (Zhang et al., 2019) 70.28 67.69 65.09 62.50 59.90 57.30 54.71 52.11 49.52 46.92
+ Ours 70.17 67.80 65.38 62.64 59.99 57.47 55.45 52.90 51.10 49.27
MART (Wang et al., 2020) 68.63 66.24 63.86 61.47 59.08 56.69 54.30 51.92 49.53 47.14
+ Ours 68.75 66.34 64.54 62.60 59.41 58.07 56.78 55.14 53.51 51.35
TRADES-AWP (Wu et al., 2020) 69.52 67.23 64.93 62.64 60.34 58.04 55.75 53.45 51.16 48.86
+ Ours 69.44 67.10 65.18 62.80 60.49 57.98 56.18 54.25 52.34 51.10
Batch Size = 100
Vanilla AT (Madry et al., 2018) 71.88 68.79 65.71 62.62 59.53 56.44 53.35 50.27 47.18 44.09
+ Ours 72.23 69.14 65.82 63.74 59.87 58.01 56.08 53.36 51.50 49.20
TRADES (Zhang et al., 2019) 70.28 67.69 65.09 62.49 59.89 57.30 54.70 52.10 49.51 46.91
+ Ours 70.36 67.79 65.22 63.58 60.62 58.31 57.04 55.51 53.38 51.58
MART (Wang et al., 2020) 68.64 66.26 63.87 61.49 59.11 56.73 54.35 51.96 49.58 47.20
+ Ours 68.75 66.35 64.50 62.61 59.50 58.02 56.77 54.97 53.51 51.40
TRADES-AWP (Wu et al., 2020) 69.52 67.22 64.92 62.62 60.31 58.01 55.71 53.41 51.11 48.81
+ Ours 69.64 67.49 64.97 63.61 61.10 59.30 57.87 56.40 54.24 52.06
Batch Size = 150
Vanilla AT (Madry et al., 2018) 71.89 68.80 65.71 62.62 59.54 56.45 53.36 50.27 47.18 44.09
+ Ours 72.06 69.17 65.96 62.86 59.68 58.92 55.09 52.39 49.90 47.97
TRADES (Zhang et al., 2019) 70.27 67.67 65.08 62.49 59.89 57.30 54.71 52.12 49.52 46.93
+ Ours 70.59 67.87 65.43 63.35 61.58 60.05 58.23 56.13 53.68 51.68
MART (Wang et al., 2020) 68.64 66.26 63.87 61.48 59.09 56.71 54.32 51.93 49.55 47.16
+ Ours 68.56 66.48 64.07 62.06 60.99 59.37 57.47 55.69 53.55 51.68
TRADES-AWP (Wu et al., 2020) 69.53 67.23 64.93 62.63 60.34 58.04 55.74 53.44 51.14 48.84
+ Ours 69.70 67.43 64.98 63.53 62.26 60.05 57.99 56.14 54.25 52.18
E.8 Compute Resources
Table 11: Training time (hours : minutes : seconds) and memory consumption (MB) for DDAD on CIFAR-10 and ImageNet-1K . This table reports the compute resources for the entire training process of DDAD (i.e., optimizing MMD + training the denoiser).
Dataset GPU Batch Size Classifier Training Time Memory
CIFAR-10 2 $\times$ NVIDIA A100 500 RN-18 00:28:17 5927
WRN-28-10 00:55:34 6276
ImageNet-1K 4 $\times$ NVIDIA A100 128 RN-50 09:52:50 97354
Table 12: Inference time (hours : minutes : seconds) for DDAD on CIFAR-10 and ImageNet-1K. This table reports the comput resources for evaluating the entire test set of CIFAR-10 (i.e., 10,000 images) and ImageNet-1K (i.e., 50,000 images).
Dataset GPU Batch Size Classifier Inference Time
CIFAR-10 1 $\times$ NVIDIA A100 100 WRN-28-10 00:00:32
ImageNet-1K 2 $\times$ NVIDIA A100 100 RN-50 00:03:08
Table 11 presents the compute resources for DDAD, which include GPU configurations, batch size, classifier, training time, and memory usage for each dataset. For CIFAR-10, using 2 NVIDIA A100 GPUs with a batch size of 500, our method’s training time is approximately 28 minutes with ResNet-18 and 55 minutes with WideResNet-28-10. The memory consumption is 5927 MB and 6276 MB, respectively. For ImageNet-1K, using 4 NVIDIA A100 GPUs with a batch size of 128, our method’s training time is approximately 10 hours, with a memory consumption of 97354 MB. Compared to AT baseline methods, DDAD offers better training efficiency (e.g., it can scale to large datasets like ImageNet-1K). This is mainly because we directly use the pre-trained classifier. Furthermore, training MMD is extremely fast (usually less than 1 minute on CIFAR-10) and we use a lightweight denoiser.
Table 12 presents the compute resources for evaluating DDAD, which include GPU configurations, batch size, classifier and inference time for each dataset. For CIFAR-10, using 1 NVIDIA A100 GPU with a batch size of 100, our method’s inference time is approximately 32 seconds over the entire test set of CIFAR-10. For ImageNet-1K, using 2 NVIDIA A100 GPUs with a batch size of 100, our method’s inference time is approximately 3 minutes over the entire test set of ImageNet-1K. Although DDAD requires training an extra denoiser and MMD-OPT, it significantly outperforms AP baselines in inference speed. Furthermore, relying on a pre-trained generative model is not always feasible, as training such models at scale can be highly resource-intensive. Therefore, considering considering the trade-off between computational cost and the performance of DDAD, we believe that training an additional detector and denoiser is feasible and worthwhile. In general, DDAD provides a more lightweight design.
Appendix F Detailed Related Work
Adversarial attacks. The discovery of adversarial examples (AEs) has raised a security concern for AI development in recent decades (Szegedy et al., 2014; Goodfellow et al., 2015). AEs are often crafted by adding imperceptible noise to clean images, which can easily mislead a classifier to make wrong predictions. The algorithms that generate AEs are called adversarial attacks. For example, the Fast Gradient Sign Method (FGSM) involves adding noise to the clean data in the direction of the gradient of the loss function with respect to the clean data (Goodfellow et al., 2015). Expanding on FGSM, the Basic Iterative Method (BIM) (Kurakin et al., 2017) iteratively applies small noises to the clean data in the direction of the gradient of the loss function, updating the input at each step to create more effective AEs than single-step methods such as FGSM. Madry et al. (2018) propose the Projected Gradient Descent (PGD), which further improves the iterative approach of BIM by adding random initialization to the input data before applying iterative gradient-based perturbations. Beyond non-targeted attacks, the Carlini & Wagner attack (C&W) specifically directs data towards a chosen target label, which crafts AEs by optimizing a specially designed objective function (Carlini & Wagner, 2017). AutoAttack (AA) (Croce & Hein, 2020a) is an ensemble of multiple adversarial attacks, which combines three non-target white-box attacks (Croce & Hein, 2020b) and one targeted black-box attack (Andriushchenko et al., 2020), which makes AA a benchmark standard for evaluating adversarial robustness. However, the computational complexity of AA is relatively high. Gao et al. (2022) propose the Minimum-margin attack (MMA), which can be used as a faster alternative to AA. Beyond computing exact gradients, Athalye et al. (2018b) propose Expectation over Transformation (EOT) to correctly compute the gradient for defenses that apply randomized transformations to the input. Athalye et al. (2018a) propose the Backward Pass Differentiable Approximation (BPDA), which approximates the gradient with an identity mapping to effectively break the defenses that leverage obfuscated gradients. According to Lee & Kim (2023), PGD+EOT is currently the best attack for denoiser-based defense methods.
Adversarial detection. The most lightweight method to defend against adversarial attacks is to detect and discard AEs in the input data. Previous studies have largely utilized statistics on hidden-layer features of deep neural networks (DNNs) to filter out AEs from test data. For example, Ma et al. (2018) utilize the local intrinsic dimensionality (LID) of DNN features as detection characteristics. Lee et al. (2018) implement a Mahalanobis distance-based score for identifying AEs. Raghuram et al. (2021) develop a meta-algorithm that extracts intermediate layer representations of DNNs, offering configurable components for detection. Deng et al. (2021) leverage a Bayesian neural network to detect AEs, which is trained by adding uniform noises to samples. Another prevalent strategy involves equipping classifiers with a rejection option. For example, Stutz et al. (2020) introduce a confidence-calibrated adversarial training framework, which guides the model to make low-confidence predictions on AEs, thereby determining which samples to reject. Similarly, Pang et al. (2022b) integrate confidence measures with a newly proposed R-Con metric to effectively separate AEs out. However, these methods, train a detector for specific classifiers or attacks, tend to neglect the modeling of data distribution, which can limit their effectiveness against unknown attacks. Recently, statistical adversarial data detection (SADD) has delivered increasing insight. For example, Gao et al. (2021) demonstrate that maximum mean discrepancy (MMD) is aware of adversarial attacks and leverage the distributional discrepancy between AEs and CEs to filter out AEs, which has been shown effective against unseen attacks. Based on this, Zhang et al. (2023) further propose a new statistic called expected perturbation score (EPS) that measures the expected score of a sample after multiple perturbations. Then, an EPS-based MMD is proposed to measure the distributional discrepancy between CEs and AEs. Despite the effectiveness of SADD, an undeniable problem of SADD-based methods is that they will discard data batches that contain AEs. To solve this problem, in this paper, we propose a new defense method that does not discard any data, while also inherits the capabilities of SADD-based detection methods.
Adversarial training. Another prominent defensive framework is adversarial training (AT). Vanilla AT (Madry et al., 2018) directly generates and incorporates AEs during the training process, forcing the model to learn the underlying distributions of AEs. Besides vanilla AT, several modifications have been developed to enhance the effectiveness of AT. For instance, at the early stage of AT, Song et al. (2019) propose to treat adversarial attacks as a domain adaptation problem and enhance the generalization of AT by minimizing the distributional discrepancy. Zhang et al. (2019) propose optimizing a surrogate loss function based on theoretical bounds. Similarly, Wang et al. (2020) explore how misclassified examples influence a model’s robustness, leading to an improved adversarial risk through regularization. From the perspective of reweighting, Ding et al. (2020) propose to reweight adversarial data with instance-dependent perturbation bounds $\epsilon$ and Zhang et al. (2021) introduce a geometry-aware instance-reweighted AT (GAIRAT) framework, which differentiates weights based on the proximity of data points to the class boundary. Wang et al. (2021) builds upon GAIRAT by leveraging probabilistic margins to reweight AEs due to their continuous nature and independence from specific perturbation paths. More recently, Zhang et al. (2024) propose to pixel-wisely reweight AEs by explicitly guiding them to focus on important pixel regions. Other modifications include improving AT using data augmentation methods (Gowal et al., 2021; Rebuffi et al., 2021) and hyper-parameter selection methods (Gowal et al., 2020; Pang et al., 2021). Although AT achieves high robustness against particular attacks, it suffers from significant degradation in clean accuracy and high computational complexity (Wong et al., 2020; Laidlaw et al., 2021; Poursaeed et al., 2021). Different from the AT framework, our method does not train a robust classifier. Instead, by directly feeding detected CEs to a pre-trained classifier, our method can effectively maintain clean accuracy. Meanwhile, by using a lightweight detector and denoiser model, our method can alleviate the computational complexity.
Denoiser-based adversarial defense. Another well-known defense framework is denoiser-based adversarial defense, which often leverages generative models to shift AEs back to their clean counterparts before feeding them into a classifier. In most literature, it is called adversarial purification (AP). Previous methods mainly focus on exploring the use of more powerful generative models for AP. For example, at the early stage of AP, Meng & Chen (2017) propose a two-step process called MagNet to remove adversarial noise by first using a detector to discard the detected AEs, and then leveraging the reconstructability of an autoencoder to purify the rest of the examples, which guides AEs towards the manifold of clean data. After MagNet, Liao et al. (2018) design a denoising UNet that can denoise AEs to their clean counterparts by reducing the distance between adversarial and clean data under high-level representations. Samangouei et al. (2018) use a GAN trained on clean examples to project AEs onto the generator’s manifold. Song et al. (2018) find that AEs lie in low-probability regions of the image distribution and propose to maximize the probability of a given test example. Naseer et al. (2020) focus on training a conditional GAN, which engages in a min-max game with a critic network, to differentiate between adversarial and clean data. Yoon et al. (2021) propose to use the denoising score-based model to purify adversarial examples. Nie et al. (2022) propose to use diffusion models to remove adversarial noise by gradually adding Gaussian noise to AEs, and then wash out the noise by solving the reverse-time stochastic differential equation. The success of recent AP methods often relies on the assumption that there will be a pre-trained generative model that can precisely estimate the probability density of the CEs (Yoon et al., 2021; Nie et al., 2022). However, even powerful generative models (e.g., diffusion models) may have an inaccurate density estimation, leading to unsatisfactory performance (Chen et al., 2024). By contrast, instead of estimating probability densities, our method directly minimizes the distributional discrepancies between AEs and CEs, leveraging the fact that identifying distributional discrepancies is simpler and more feasible than estimating density. Nayak et al. (2023) propose to use MMD as a regularizer during the training of the denoiser. Different from their work, we use an optimized version of MMD (i.e., MMD-OPT), which is more sensitive to adversarial attacks. Furthermore, the MMD-OPT serves not only as a ‘guider’ during training to help minimize the distributional discrepancy between AEs and CEs, but also a ‘detector’ that helps distinguish AEs and CEs.
Appendix G Discussions on Batch-wise Detections
Benefits of using batch-wise statistical hypothesis test. A main benefit of using a batch-wise statistical hypothesis test is that it can effectively control the false alarm rate. For example, for DDAD, we set the maximum false alarm rate to be 5%. Fang et al. (2022) theoretically prove that for instance-wise detection methods to work perfectly, there must be a gap in the support set between IID and out-of-distribution (OOD) data. This theory also applies to adversarial problems, but such a support set does not exist in adversarial settings, making perfect instance-wise detection generally difficult.
Limitation and its solutions for user inference. DDAD leverages statistics based on distributional discrepancies (i.e., MMD-OPT), which requires the data to be processed in batches for adversarial detection. However, when the batch size is too small, the stability of DDAD will be affected (see Figure 2). To address this issue, for user inference, single samples provided by the user can be dynamically stored in a queue. Once the queue accumulates enough samples to form a batch, our method can then process the batch collectively using the proposed approach. A direct cost of this solution is the waiting time, as the system must accumulate enough samples (e.g., 50 samples) to form a batch before processing. However, in scenarios where data arrives quickly, the waiting time is typically very short, making this approach feasible for many real-time applications. For applications with stricter latency requirements, the batch size can be dynamically adjusted based on the incoming data rate to minimize waiting time. For instance, if the system detects a lower data arrival rate, it can process smaller batches to ensure timely responses. Overall, it is a trade-off problem: using our method for user inference can obtain high robustness, but the cost is to wait for batch processing. Based on the performance improvements our method obtains over the baseline methods, we believe the cost is feasible and acceptable. Another possible solution is to find more robust statistics that can measure distributional discrepancies with fewer samples. Recently, measuring the expected score of a sample after multiple perturbations has proven useful for this purpose (Zhang et al., 2023). However, computing the expected score is time-consuming. We emphasize that this paper primarily focuses on the relationship between distributional discrepancies and adversarial risk, aiming to inspire the design of a new defense method. We leave it as future work.
Practicability beyond user inference. On the other hand, our method is not necessarily used for user inference. Instead, our method is suitable for cleaning the data before fine-tuning the underlying model. In many domains, obtaining large quantities of high-quality data is challenging due to factors such as cost, privacy concerns, or the rarity of specific data. As a result, all possible samples with clean information are critical in these data-scarce domains. Then, a practical scenario is that there exists a pre-trained model on a large-scale dataset (e.g., a DNN trained on ImageNet-1K) and clients want to fine-tune the model to perform well on downstream tasks. If the data for downstream tasks contain AEs, our method can be applied to batch-wisely clean the data before fine-tuning the underlying model.